solve the system using the linear combination method: 5x+y=6 -5x+3y=-22
Answer by funmath(2933)

I didn't answer this yesterday because I wasn't positive that combination was the same thing as elimination or addition or subtraction methods, they call it so many things. However, since this is still here, I will tell you what I think you want. If I'm using the wrong method, I appologize.
:
Solve the system using the linear combination method:
Add the two equations together and the x's are eliminated.
5x+y=6
-5x+3y=-22
_________
4y=-16
4y/4=-16/4
y=-4
Substitute y=-4 into one of the original equations and solve for x:
5x+(-4)=6
5x-6=6
5x-6+6=6+6
5x=10
5x/5=10/5
x=2
The solution is (x,y)=(2,-4)
We can check our answers to be sure we're right by substituting (2,-4) into both equations and see if they both equal.
5(2)+(-4)=6
10-4=6
6=6 First equation checks out.
-5(2)+3(-4)=-22
-10-12=-22
-22=-22 Second equation checks out too. We're right!!!
Happy Calculating!!!